OpenStudy (anonymous):

solve the proportion. 2x+2/4=x+1/6

OpenStudy (anonymous):

Multiply top left times bottom right... that equals bottom left times top right.

OpenStudy (anonymous):

(2x + 2)(6) = 4(x + 1 ) distribute

OpenStudy (anonymous):

12x + 12 = 4x + 4

OpenStudy (anonymous):

Subtract 4x 8x + 12 = 4 subtract 12 8x = -8 divide by 8 x = -1

OpenStudy (anonymous):

would it be -1 or 1

OpenStudy (anonymous):

4 - 12 = -8 so 8x = -8 x = -1